OpenStudy (fools101):

1 - Zeros of Polynomials Zeros of Polynomial Functions 1. Recreation: The distance d, in centimeters, that a diving board bends below its resting position when you stand at its end is dependent on your distance x in meters, from the stabilized point. This relationship can be modeled by the function: d(x) = -4x3 + x2 Evaluate d(x) for x = 2 meters from the stabilized point.

OpenStudy (fools101):

please help!

OpenStudy (anonymous):

d(2) = -4 * 8 + 4 =-28

jimthompson5910 (jim_thompson5910):

d(x) = -4x^3 + x^2 d(2) = -4(2)^3 + (2)^2 d(2) = -4(8) + 4 d(2) = -32 + 4 d(2) = -28

OpenStudy (fools101):

Thank u!!! JIM!!

jimthompson5910 (jim_thompson5910):

yw